OpenStudy (anonymous):

A man wants to set aside $3600 at 8% interest compounded quarterly. Find the amount of money at the end of 3 years.

OpenStudy (anonymous):

\[3600\times\left(1+\frac{.08}{4}\right)^{12}\] and a calculator

OpenStudy (anonymous):

I did that the first time. $4,565.67 so I was right the first time? Quarterly threw me off a bit.

OpenStudy (austinl):

\[ A = P(1+\frac{r}{n})^{nt}\] P = Principal Amount r = annual rate of interest t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year

OpenStudy (austinl):

Example: An amount of $1,500.00 is deposited in a bank paying an annual interest rate of 4.3%, compounded quarterly. What is the balance after 6 years? Solution: Using the compound interest formula, we have that P = 1500, r = 4.3/100 = 0.043, n = 4, t = 6. Therefore, \[A = $1,500(1+\frac{0.043}{4})^{4(6)}\approx $1,938.84\] So, the balance after 6 years is approximately $1,938.84.

OpenStudy (anonymous):

Rate for per year=8% Rate per quarter=8/4%=2% No. of quarters 3*4=12 \[A=3600\left( 1+\frac{ 2 }{100 } \right)^{12}\] calculate